OpenStudy (anonymous):

A bacteria culture begins with 15 bacteria which double in amount every hour. How many bacteria exist in the culture at the 13th hour? 61,440 782,580 7,971,615 23,914,845

OpenStudy (anonymous):

@D3xt3R @FriedRice

OpenStudy (anonymous):

\[P(t)=B*2^{t-1}\]\[P(t)=15*2^{t-1}\]\[P(13)=15*2^{13-1}\]\[P(13)=15*2^{12}\]\[\boxed{P(13)=61440}\]

OpenStudy (anonymous):

thank you so much